OpenStudy (anonymous):

Rewrite in vertex form y=x^2+14x+50

OpenStudy (anonymous):

x = 7. my guess but idk

OpenStudy (anonymous):

idk really know but, that just a guess :)

OpenStudy (anonymous):

Vertex: (−7,1)

OpenStudy (anonymous):

let me explain the process.

OpenStudy (anonymous):

The general form of the quadratic function is f(x) =ax2 + bx + c. To convert the standard form(y = ax2 + bx + c) of a function into vertex form(y = a(x - h)2 + k), we have to write the equation in the complete square form and vertex(h, k) is given by: h = −b2a k = c - b24a just gotta plug it in and you get (-7,1)

OpenStudy (anonymous):

since There is no A you would use one, since there is always a 1 in front of something when you don't see it

OpenStudy (anonymous):

Ok so when i used the formula \[y=a(x-h)^2+k\] I got \[y=14(x-x)^2+50\]

OpenStudy (anonymous):

is that correct when rewriting it in vertex form?

OpenStudy (anonymous):

@cantthink96

OpenStudy (anonymous):

i'm ever so sorry i didn't realize thats what the question was

OpenStudy (anonymous):

you're going to need to find h and k before you can put it in vertex fom

OpenStudy (anonymous):

Yeh so did I do it right by rewriting it that way?

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

h(x-coordinate) = −b2a k(y - coordinate) = c - b24a

OpenStudy (anonymous):

Ok thank you for the help!

OpenStudy (anonymous):

hold on its -b/2a and c- b^2/4a